A 1.00-g cork ball with charge 2.00 *C is

suspended vertically on a 0.500-m-long light string in the

presence of a uniform, downward-directed electric field of

magnitude E ! 1.00 & 105 N/C. If the ball is displaced

slightly from the vertical, it oscillates like a simple pendulum.

(a) Determine the period of this oscillation.

(b) Should gravity be included in the calculation for part

(a)? Explain.


1
Expert's answer
2022-03-20T18:47:53-0400

"T=2\pi\sqrt{\frac la}=2\pi\sqrt{\frac{ml}{mg+F}}=2\pi\sqrt{\frac{ml}{mg+qE}}\approx \sqrt{\frac{ml}{qE}}=5~s."

"mg\ll qE," gravity won't be included in the calculation.